# Sharpness of Minima in Deep Matrix Factorization: Exact Expressions

### v4

@@ 2 Notation, Preliminaries, and Problem Setup
$\lim_{\Delta t\rightarrow 0}\frac{f(\bm{X}+(\Delta t+t)\bm{U})-f(\bm{X}+t\bm{U})}{\Delta t}\Biggl|_{t=0}=\frac{\partial f(\bm{X}+t\bm{U})}{\partial t}\Biggl|_{t=0}.$ (10)
This is known as the Gâteaux derivative, which represents the change in $f(\bm{X})$ under a perturbation in the direction of $\bm{U}$. By the same reasoning, we obtain the following result.
Lemma 2.
The second directional derivative of $f$ at $\bm{X}$ in the direction $\bm{U}\in\mathbb{R}^{K\times L}$ is given by
$D_{\bm{U}}^{2}f(\bm{X})=\frac{\partial^{2}}{\partial t^{2}}f(\bm{X}+t\bm{U})\Big|_{t=0}.$ (11)
@@
Directional Second Derivatives and Maximum Eigenvalue. Consider the following objective function for our real-valued matrix-variable function $f$.
$f(\bm{W}_{1},\bm{W}_{2},\dots,\bm{W}_{L})=\norm{\bm{M}-\bm{W}_{L}\bm{W}_{L-1}\cdots\bm{W}_{1}}_{F}^{2},$ (12)
where $\bm{W}_{i}\in\mathbb{R}^{d_{i}\times d_{i-1}}$ and $\bm{M}\in\mathbb{R}^{d_{L}\times d_{0}}$ for all $i\in[L]$. In this setting, we can define the largest eigenvalue of the $\nabla^{2}f(\bm{W}_{1},\bm{W}_{2},\dots,\bm{W}_{L})$ at an arbitrary point in the parameter space as follows:
$\lambda_{\max}(\nabla^{2}f(\bm{W}_{1},\bm{W}_{2},\dots,\bm{W}_{L}))=\max_{\begin{subarray}{c}\bm{U}_{1},\bm{U}_{2},\cdots,\bm{U}_{L}:\\ \sum_{i=1}^{L}\norm{\bm{U}_{i}}_{F}^{2}=1\end{subarray}}\frac{d^{2}}{dt^{2}}f(\bm{W}_{1}+t\bm{U}_{1},\cdots,\bm{W}_{L}+t\bm{U}_{L})\Bigr|_{t=0}.$ (13)
@@
Lemma 3.
For any $\bm{W}^{*}_{1},\bm{W}^{*}_{2},\cdots,\bm{W}^{*}_{L}]$ such that $\bm{M}=\prod_{j=1}^{L}\bm{W}^{*}_{j}$, the directional second derivative is given by
$\nabla^{2}f(\bm{W}^{*}_{1},\cdots,\bm{W}^{*}_{L})[\bm{U}_{1},\cdots,\bm{U}_{L}]=2\norm{\sum_{i=1}^{L}\Biggl[\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\bm{U}_{i}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)\Biggr]}_{F}^{2}.$ (14)
@@
We study the sharpness of the loss landscape near any global minimum in deep matrix factorization problems. We consider the following optimization problem
$\min_{\bm{w}\in\mathbb{R}^{N}}\mathcal{L}(\bm{w}):=\norm{\bm{M}-\bm{W}_{L}\bm{W}_{L-1}\cdots\bm{W}_{1}}_{F}^{2},$ (15)
where $\bm{w}=\operatorname{vec}([\bm{W}_{1},\bm{W}_{2},\dots,\bm{W}_{L}])$ denotes the collection of all parameters, and
@@
is the total number of parameters in the model. $\bm{M}\in\mathbb{R}^{d_{L}\times d_{0}}$ denotes the matrix contains the parameters subject to factorization, $L\geq 2$ denotes the depth of factorization and $\bm{W}_{i}\in\mathbb{R}^{d_{i}\times d_{i-1}}$ is the $i^{th}$ factor (layer). This objective is analogous to that of deep linear neural networks. To guarantee the feasibility of factorization at all points in $\mathbb{R}^{d_{L}\times d_{0}}$, we require
@@
$\operatorname{rank}(\bm{W}_{L}\bm{W}_{L-1}\cdots\bm{W}_{1})\leq\operatorname{min}\{\operatorname{rank}(\bm{W}_{1}),\operatorname{rank}(\bm{W}_{2}),\dots,\operatorname{rank}(\bm{W}_{L})\}.$ (18)
Define the set of global minima of $\mathcal{L}(\bm{w)}$ as
$\Omega:=\operatorname*{arg\,min}_{\bm{w}\in\mathbb{R}^{N}}\mathcal{L}(\bm{w})=\left\{\bm{w}\in\mathbb{R}^{N}:\prod_{i=1}^{L}\bm{W}_{i}=\bm{M}\right\}.$ (19)
Local Minima. Laurent & Brecht (2018) showed that all local minima of deep linear networks with convex and differentiable loss are global if the layers satisfy (17), i.e., hidden layers are at least as wide as input and output layers. In particular, they proved a more general theorem concerning real-valued functions that take as input a product of matrices. Thus, a corollary of Laurent & Brecht (2018, Theorem 1) is that overparameterized deep matrix factorization does not have spurious minima, i.e., all local minima are global.
3 Deep Overparameterized Scalar Factorization
Before we delve into our general results, we first investigate the deep overparameterized scalar factorization, i.e., a special case of deep matrix factorization in which the first and last layers are vectors. This simplified problem setup reveals the key proof techniques used to prove our general result in Section 4.
Theorem 4.
@@ 3 Deep Overparameterized Scalar Factorization
$\mathcal{L}(\bm{w}):=({m-\bm{w}_{L}\bm{W}_{L-1}\cdots\bm{W}_{2}\bm{w}_{1}})^{2},$ (20)
where $m\in\mathbb{R}$, $d_{0}=d_{L}=1$, $\bm{w}_{L}\in\mathbb{R}^{1\times d_{L-1}}$ and $\bm{w}_{1}\in\mathbb{R}^{d_{1}\times 1}$. For hidden factors (layers), i.e, for all $i\in\{2,3,\cdots,L-1\}$, we have $\bm{W}_{i}\in\mathbb{R}^{d_{i}\times d_{i-1}}$. Then, $\forall\bm{w}^{*}\in\Omega$,
$\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w^{*}}))=2\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\bm{W}_{j}^{*}\Bigr)^{2}\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)^{2}.$ (21)
The proof appears in Appendix B.
4 Overparameterized Deep Matrix Factorization
We now consider the general deep matrix factorization problem. In this section, we prove our main result, which is closed-form expression for the maximum eigenvalue of the Hessian for any global minimum to the objective (15).
Theorem 5.
If $\bm{w}^{*}\in\Omega$ then
$\lambda_{\max}\bigl(\nabla^{2}\mathcal{L}(\bm{w}^{*})\bigr)=2\sigma_{\max}\Bigl(\sum_{i=1}^{L}\bm{B}_{i}^{\top}\bm{B}_{i}\otimes\bm{A}_{i}\bm{A}_{i}^{\top}\Bigr),$ (22)
where $\bm{A}_{k}=\prod_{i=k+1}^{L}\bm{W}_{i}^{*}$ and $\bm{B}_{k}=\prod_{i=1}^{k-1}\bm{W}_{i}^{*}$.
The proof appears in Appendix C.1. Note that the deep overparameterized scalar factorization is a special case of deep matrix factorization where both $\bm{B}_{i}^{\top}\bm{B}_{i}$ and $\bm{A}_{i}\bm{A}_{i}^{\top}$ reduce to scalars. In that special case, we recover Theorem 21. Another corollary of Theorem 5 is the maximum Hessian eigenvalue for the classical (depth-$2$) matrix factorization problem. This result may be of independent interest as the expression simplifies considerably.
Corollary 6.
@@ 4 Overparameterized Deep Matrix Factorization
$\mathcal{L}(\bm{L},\bm{R})=\norm{\bm{M}-\bm{L}\bm{R}^{\top}}_{F}^{2},$ (23)
where $\bm{M}\in\mathbb{R}^{d_{L}\times d_{0}}$ is the target matrix and $\bm{L}\in\mathbb{R}^{d_{L}\times k}$, $\bm{R}\in\mathbb{R}^{d_{0}\times k}$. To ensure the feasibility of the factorization every point in $\mathbb{R}^{d_{L}\times d_{0}}$, we choose $k\geq\min\{d_{0},d_{L}\}$. We define the set of minimizers as follows:
$\Omega:=\operatorname*{arg\,min}_{\bm{L},\bm{R}}\mathcal{L}(\bm{L},\bm{R})=\left\{(\bm{L},\bm{R}):\bm{M}=\bm{L}\bm{R}^{T}\right\}.$ (24)
If $(\bm{L},\bm{R})\in\Omega$ then
$\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{L},\bm{R}))=2(\sigma_{\max}(\bm{L})^{2}+\sigma_{\max}(\bm{R})^{2}).$ (25)
@@
We have from Theorem 5 that
$\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{L},\bm{R}))=2\sigma_{\max}(\bm{I}\otimes\bm{L}\bm{L}^{\top}+\bm{R}\bm{R}^{\top}\otimes\bm{I}).$ (26)
Using the fact from Horn & Johnson (1994, Theorem 4.4.5), we can write
$2\sigma_{\max}(\bm{I}\otimes\bm{L}\bm{L}^{\top}+\bm{R}\bm{R}^{\top}\otimes\bm{I})=2\sigma_{\max}(\bm{I}\otimes\bm{L}\bm{L}^{\top})+2\sigma_{\max}(\bm{R}\bm{R}^{\top}\otimes\bm{I}).$ (27)
Note that for any matrix $\bm{A}$ and $\bm{B}$, $\sigma_{\max}(\bm{A}\otimes\bm{B})=\sigma_{\max}(\bm{A})\sigma_{\max}(\bm{B})$, and $\sigma_{\max}(\bm{A}\bm{A}^{\top})=\sigma_{\max}(\bm{A}^{\top}\bm{A})=\sigma_{\max}(\bm{A})^{2}$. Hence,
$2(\sigma_{\max}(\bm{I}\otimes\bm{L}\bm{L}^{\top})+\sigma_{\max}(\bm{R}\bm{R}^{\top}\otimes\bm{I}))=2(\sigma_{\max}(\bm{L})^{2}+\sigma_{\max}(\bm{R})^{2}).$ (28)
@@ 5 Flatness
In this section, we reveal remarkable aspects of the loss landscape of general deep matrix factorization problems. First, we show that an optimal solution in deep overparameterized scalar factorization problem is flat if and only if the product of spectral norms of left and right intermediate networks is constant across layers. This is a direct consequence of Theorem 21.
Corollary 7.
@@
$\mathcal{L}(\bm{w}):=({m-\bm{w}_{L}\bm{W}_{L-1}\cdots\bm{W}_{2}\bm{w}_{1}})^{2},$ (29)
where $m\in\mathbb{R}$, $d_{0}=d_{L}=1$, $\bm{w}_{L}\in\mathbb{R}^{1\times d_{L-1}}$ and $\bm{w}_{1}\in\mathbb{R}^{d_{1}\times 1}$. For hidden factors (layers), i.e, for all $i\in\{2,3,\cdots,L-1\}$, we have $\bm{W}_{i}\in\mathbb{R}^{d_{i}\times d_{i-1}}$. Then, $\bm{w}^{*}\in\Omega_{F}$ if and only if
$\sigma_{\max}(\bm{A}_{k})\sigma_{\max}(\bm{B}_{k})=|m|^{1-\frac{1}{L}}\quad\forall k\in[L],$ (30)
where $\bm{A}_{k}=\prod_{i=k+1}^{L}\bm{W}_{i}^{*}$ and $\bm{B}_{k}=\prod_{i=1}^{k-1}\bm{W}_{i}^{*}$.
@@
For a minimizer $\bm{w}^{*}$, let us assume that $\sigma_{\max}(\bm{A}_{k})\sigma_{\max}(\bm{B}_{k})=|m|^{1-\frac{1}{L}}$ for all $k\in[L]$. Then, by Theorem 21, $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))=2L\times m^{2(1-1/L)}$. This implies that $\bm{w}^{*}$ is flat (Mulayoff & Michaeli, 2020, Theorem 1). Now, assume that $\bm{w}^{*}$ is flat. (Mulayoff & Michaeli, 2020, Theorem 2) showed that for any flat minimum $\bm{w}^{*}$,
$\sigma_{\max}(\bm{A}_{k})=|m|^{1-\frac{k}{L}}\quad\text{and}\quad\sigma_{\max}(\bm{B}_{k})=|m|^{\frac{k-1}{L}}\quad\forall k\in[L].$ (31)
Hence, $\sigma_{\max}(\bm{A}_{k})\sigma_{\max}(\bm{B}_{k})=|m|^{1-\frac{1}{L}}$ for all $k\in[L]$. ∎
Second, we show that being flat in depth-2 matrix factorization is equivalent to being spectral-norm balanced. This result is a direct consequence of Corollary 25 and AM-GM inequality (Josz, 2025, see also Lemma 5).
Corollary 8.
@@
$\mathcal{L}(\bm{L},\bm{R})=\norm{\bm{M}-\bm{L}\bm{R}^{\top}}_{F}^{2},$ (32)
where $\bm{M}\in\mathbb{R}^{d_{L}\times d_{0}}$ is the target matrix and $\bm{L}\in\mathbb{R}^{d_{L}\times k}$, $\bm{R}\in\mathbb{R}^{d_{0}\times k}$ such that $k\geq\min\{d_{0},d_{L}\}$. Then, $(\bm{L}^{*},\bm{R}^{*})\in\Omega_{F}$ if and only if
$\sigma_{\max}(\bm{L}^{*})=\sigma_{\max}(\bm{R}^{*})=\sqrt{\sigma_{\max}(\bm{M})}.$ (33)
Remark 9.
We discuss further implications of this equivalence in the subsequent section.
@@
Corollary 10.
Consider the optimization objective in (15). If $\bm{w}^{*}\in\Omega$ then
$\lambda_{\max}\bigl(\nabla^{2}\mathcal{L}(\bm{w}^{*})\bigr)\leq 2\sum_{k=1}^{L}\sigma_{\max}(\bm{A}_{k})^{2}\sigma_{\max}(\bm{B}_{k})^{2},$ (34)
where $\bm{A}_{k}=\prod_{i=k+1}^{L}\bm{W}_{i}^{*}$ and $\bm{B}_{k}=\prod_{i=1}^{k-1}\bm{W}_{i}^{*}$ (see Appendix B). This implies that if $\bm{w}^{*}$ satisfies $\sigma_{\max}(\bm{A}_{k})\sigma_{\max}(\bm{B}_{k})=\sigma_{\max}(\bm{M})^{1-\frac{1}{L}}$ for all $k\in[L]$, then $\bm{w}^{*}$ is flat.
@@
Mulayoff & Michaeli (2020) showed that for any $\bm{w}^{*}\in\Omega$,
$\lambda_{\max}\bigl(\nabla^{2}\mathcal{L}(\bm{w}^{*})\bigr)\geq 2L\sigma_{\max}(\bm{M})^{2(1-1/L)}.$ (35)
Now, assume that $\sigma_{\max}(\bm{A}_{k})\sigma_{\max}(\bm{B}_{k})=\sigma_{\max}(\bm{M})^{1-\frac{1}{L}}$ for all $k\in[L]$. Then, by (34),
$\lambda_{\max}\bigl(\nabla^{2}\mathcal{L}(\bm{w}^{*})\bigr)=2L\sigma_{\max}(\bm{M})^{2(1-1/L)}.$ (36)
This implies that $\bm{w}^{*}$ is flat (Mulayoff & Michaeli, 2020, Theorem 1). ∎

### v5

@@ 2.1 Gâteaux Derivatives
$\lim_{\Delta t\rightarrow 0}\frac{f\left(\mathbf{X}+\left(\Delta t+t\right)\mathbf{U}\right)-f\left(\mathbf{X}+t\mathbf{U}\right)}{\Delta t}\Biggl|_{t=0}.$ (12)
This expression can be rewritten as
$\frac{\partial f\left(\mathbf{X}+t\mathbf{U}\right)}{\partial t}\Biggl|_{t=0}.$ (13)
This is known as the Gâteaux derivative, which represents the change in $f\left(\mathbf{X}\right)$ under a perturbation in the direction of $\mathbf{U}$. By the same reasoning, we obtain the following result.
Lemma 2.1.
The second directional derivative of $f$ at $\mathbf{X}$ in the direction $\mathbf{U}\in\mathbb{R}^{K\times L}$ is given by
$D_{\mathbf{U}}^{2}f\left(\mathbf{X}\right)=\frac{\partial^{2}}{\partial t^{2}}f\left(\mathbf{X}+t\mathbf{U}\right)\Big|_{t=0}.$ (14)
@@
2.2 Directional Second Derivatives and Maximum Eigenvalue
Consider the following objective function for our real-valued matrix-variable function $f$.
$f\left(\mathbf{W}_{1},\mathbf{W}_{2},\dots,\mathbf{W}_{L}\right)=\norm{\mathbf{M}-\mathbf{W}_{L}\mathbf{W}_{L-1}\cdots\mathbf{W}_{1}}_{F}^{2},$ (15)
where $\mathbf{W}_{i}\in\mathbb{R}^{d_{i}\times d_{i-1}}$ and $\mathbf{M}\in\mathbb{R}^{d_{L}\times d_{0}}$ for all $i\in[L]$. In this setting, we define the maximum eigenvalue of the $\nabla^{2}f\left(\mathbf{W}_{1},\mathbf{W}_{2},\dots,\mathbf{W}_{L}\right)$ at an arbitrary point in the parameter space as follows:
$\lambda_{\max}\left(\nabla^{2}f\left(\mathbf{W}_{1},\mathbf{W}_{2},\dots,\mathbf{W}_{L}\right)\right)=$
$\max_{\begin{subarray}{c}\mathbf{U}_{1},\mathbf{U}_{2},\cdots,\mathbf{U}_{L}:\\ \sum_{i=1}^{L}\norm{\mathbf{U}_{i}}_{F}^{2}=1\end{subarray}}\frac{d^{2}}{dt^{2}}f\left(\mathbf{W}_{1}+t\mathbf{U}_{1},\cdots,\mathbf{W}_{L}+t\mathbf{U}_{L}\right)\Bigr|_{t=0}.$ (16)
@@ 2.2 Directional Second Derivatives and Maximum Eigenvalue
Lemma 2.2.
For any $\left[\mathbf{W}^{*}_{1},\mathbf{W}^{*}_{2},\cdots,\mathbf{W}^{*}_{L}\right]$ such that $\mathbf{M}=\prod_{j=1}^{L}\mathbf{W}^{*}_{j}$, the directional second derivative along $\mathbf{U}_{1},\cdots,\mathbf{U}_{L}]$ is given by
$\nabla^{2}f\left(\mathbf{W}^{*}_{1},\cdots,\mathbf{W}^{*}_{L}\right)[\mathbf{U}_{1},\cdots,\mathbf{U}_{L}]=$
$2\norm{\sum_{i=1}^{L}\left[\left(\prod_{j=i+1}^{L}\mathbf{W}^{*}_{j}\right)\mathbf{U}_{i}\left(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\right)\right]}_{F}^{2}.$ (17)
@@
In this paper, we study the sharpness of the loss landscape of deep matrix factorization problems near any global minimizer. We consider the following optimization problem
$\min_{\mathbf{w}\in\mathbb{R}^{N}}\mathcal{L}\left(\mathbf{w}\right):=\norm{\mathbf{M}-\mathbf{W}_{L}\mathbf{W}_{L-1}\cdots\mathbf{W}_{1}}_{F}^{2},$ (18)
where $\mathbf{w}=\operatorname{vec}\left([\mathbf{W}_{1},\mathbf{W}_{2},\dots,\mathbf{W}_{L}]\right)$ denotes the collection of all parameters, and
@@
is the total number of parameters in the model. $\mathbf{M}\in\mathbb{R}^{d_{L}\times d_{0}}$ denotes the matrix that contains the parameters subject to factorization, $L\geq 2$ denotes the depth of factorization and $\mathbf{W}_{i}\in\mathbb{R}^{d_{i}\times d_{i-1}}$ is the $i$th factor (layer). This objective is analogous to that of deep linear neural networks. To guarantee the feasibility of factorization for all points in $\mathbb{R}^{d_{L}\times d_{0}}$, we require
@@
$\operatorname{rank}\left(\mathbf{W}_{L}\cdots\mathbf{W}_{1}\right)\leq\min\left\{\operatorname{rank}\left(\mathbf{W}_{1}\right),\dots,\operatorname{rank}\left(\mathbf{W}_{L}\right)\right\}.$ (21)
We define the set of global minima of $\mathcal{L}(\mathbf{w})$ as
$\Omega:=\operatorname*{arg\,min}_{\mathbf{w}\in\mathbb{R}^{N}}\mathcal{L}(\mathbf{w})=\left\{\mathbf{w}\in\mathbb{R}^{N}:\prod_{i=1}^{L}\mathbf{W}_{i}=\mathbf{M}\right\}.$ (22)
2.3 Benign Landscape
Laurent and Brecht (2018) showed that all local minima of deep linear networks with convex and differentiable loss are global if the layers satisfy (20), i.e., hidden layers are at least as wide as input and output layers. In particular, they proved a more general theorem concerning real-valued functions that take as input a product of matrices. Thus, a corollary of Laurent and Brecht (2018, Theorem 1) is that overparameterized deep matrix factorization does not have spurious minima, i.e., all local minima are global. Thus, in this paper, we can explicitly focus on global minima, without loss of generality.
3 Overparameterized Deep Matrix Factorization
We now consider the general deep matrix factorization problem. In this section, we prove our main result, which is a closed-form expression for the maximum eigenvalue of the Hessian for any global minimum to the objective (18).
Theorem 3.1.
If $\mathbf{w}^{*}\in\Omega$ then
$\lambda_{\max}\left(\nabla^{2}\mathcal{L}\left(\mathbf{w}^{*}\right)\right)=2\sigma_{\max}\left(\sum_{i=1}^{L}\mathbf{B}_{i}^{\top}\mathbf{B}_{i}\otimes\mathbf{A}_{i}\mathbf{A}_{i}^{\top}\right),$ (23)
where $\mathbf{A}_{k}=\prod_{i=k+1}^{L}\mathbf{W}_{i}^{*}$ and $\mathbf{B}_{k}=\prod_{i=1}^{k-1}\mathbf{W}_{i}^{*}$.
The proof appears in Appendix C.1. Note that the deep overparameterized scalar factorization is a special case of deep matrix factorization where both $\mathbf{B}_{i}^{\top}\mathbf{B}_{i}$ and $\mathbf{A}_{i}\mathbf{A}_{i}^{\top}$ reduce to scalars. In that special case, we have Corollary 3.2 below. Another corollary of Theorem 3.1 is the maximum Hessian eigenvalue for the classical (depth-$2$) matrix factorization problem. This result may be of independent interest as the expression simplifies considerably.
Corollary 3.2.
@@ 3 Overparameterized Deep Matrix Factorization
$\mathcal{L}(\mathbf{w}):=({m-\mathbf{w}_{L}\mathbf{W}_{L-1}\cdots\mathbf{W}_{2}\mathbf{w}_{1}})^{2},$ (24)
where $m\in\mathbb{R}$, $d_{0}=d_{L}=1$, $\mathbf{w}_{L}\in\mathbb{R}^{1\times d_{L-1}}$ and $\mathbf{w}_{1}\in\mathbb{R}^{d_{1}\times 1}$. For hidden factors (layers), i.e., for all $i\in\{2,3,\cdots,L-1\}$, we have $\mathbf{W}_{i}\in\mathbb{R}^{d_{i}\times d_{i-1}}$. Then, for all $\mathbf{w}^{*}\in\Omega$,
$\lambda_{\max}\left(\nabla^{2}\mathcal{L}\left(\mathbf{w}^{*}\right)\right)=$
$2\sum_{i=1}^{L}\sigma_{\max}\left(\prod_{j=i+1}^{L}\mathbf{W}_{j}^{*}\right)^{2}\sigma_{\max}\left(\prod_{j=1}^{i-1}\mathbf{W}_{j}^{*}\right)^{2}.$ (25)
We also provide a self-contained proof of this corollary in Appendix B.
Corollary 3.3.
@@
$\mathcal{L}(\mathbf{L},\mathbf{R})=\norm{\mathbf{M}-\mathbf{L}\mathbf{R}^{\top}}_{F}^{2},$ (26)
where $\mathbf{M}\in\mathbb{R}^{d_{L}\times d_{0}}$ is the target matrix and $\mathbf{L}\in\mathbb{R}^{d_{L}\times k}$, $\mathbf{R}\in\mathbb{R}^{d_{0}\times k}$. To ensure the feasibility of the factorization for every point in $\mathbb{R}^{d_{L}\times d_{0}}$, we choose $k\geq\min\{d_{0},d_{L}\}$. We define the set of minimizers as follows:
$\Omega:=\operatorname*{arg\,min}_{\mathbf{L},\mathbf{R}}\mathcal{L}\left(\mathbf{L},\mathbf{R}\right)=\left\{\left(\mathbf{L},\mathbf{R}\right):\mathbf{M}=\mathbf{L}\mathbf{R}^{\top}\right\}.$ (27)
If $(\mathbf{L},\mathbf{R})\in\Omega$, then
$\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{L},\mathbf{R}))=2(\sigma_{\max}(\mathbf{L})^{2}+\sigma_{\max}(\mathbf{R})^{2}).$ (28)
@@
We have from Theorem 3.1 that $\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{L},\mathbf{R}))$
$=2\sigma_{\max}\left(\mathbf{I}\otimes\mathbf{L}\mathbf{L}^{\top}+\mathbf{R}\mathbf{R}^{\top}\otimes\mathbf{I}\right)$ (29)
$=2\left(\sigma_{\max}\left(\mathbf{I}\otimes\mathbf{L}\mathbf{L}^{\top}\right)+\sigma_{\max}\left(\mathbf{R}\mathbf{R}^{\top}\otimes\mathbf{I}\right)\right)$ (30)
$=2\left(\sigma_{\max}\left(\mathbf{L}\right)^{2}+\sigma_{\max}\left(\mathbf{R}\right)^{2}\right)$ (31)
Using the fact from Horn and Johnson (1994, Theorem 4.4.5), we can obtain (30) from (29). Note that for any matrix $\mathbf{A}$ and $\mathbf{B}$, $\sigma_{\max}(\mathbf{A}\otimes\mathbf{B})=\sigma_{\max}(\mathbf{A})\sigma_{\max}(\mathbf{B})$, and $\sigma_{\max}(\mathbf{A}\mathbf{A}^{\top})=\sigma_{\max}(\mathbf{A}^{\top}\mathbf{A})=\sigma_{\max}(\mathbf{A})^{2}$. Hence, we rewrite (30) as (31). ∎
@@ 4 Flatness
In this section, we reveal remarkable aspects of the loss landscape of general deep matrix factorization problems. First, we show that an optimal solution of the deep overparameterized scalar factorization problem is flat if and only if the product of spectral norms of left and right intermediate networks is constant across layers. This is a direct consequence of Corollary 3.2.
Corollary 4.1.
@@
$\mathcal{L}\left(\mathbf{w}\right):=({m-\mathbf{w}_{L}\mathbf{W}_{L-1}\cdots\mathbf{W}_{2}\mathbf{w}_{1}})^{2},$ (32)
where $m\in\mathbb{R}$, $d_{0}=d_{L}=1$, $\mathbf{w}_{L}\in\mathbb{R}^{1\times d_{L-1}}$ and $\mathbf{w}_{1}\in\mathbb{R}^{d_{1}\times 1}$. For hidden factors (layers), i.e., for all $i\in\{2,3,\cdots,L-1\}$, we have $\mathbf{W}_{i}\in\mathbb{R}^{d_{i}\times d_{i-1}}$. Then, $\mathbf{w}^{*}\in\Omega_{F}$ if and only if
$\sigma_{\max}(\mathbf{A}_{k})\sigma_{\max}(\mathbf{B}_{k})=|m|^{1-\frac{1}{L}}\quad\forall k\in[L],$ (33)
where $\mathbf{A}_{k}=\prod_{i=k+1}^{L}\mathbf{W}_{i}^{*}$ and $\mathbf{B}_{k}=\prod_{i=1}^{k-1}\mathbf{W}_{i}^{*}$.
@@
For a minimizer $\mathbf{w}^{*}$, let us assume that $\sigma_{\max}(\mathbf{A}_{k})\sigma_{\max}(\mathbf{B}_{k})=|m|^{1-\frac{1}{L}}$ for all $k\in[L]$. Then, by Corollary 3.2, $\lambda_{\max}(\nabla^{2}\mathcal{L}(\mathbf{w}^{*}))=2L\times m^{2(1-1/L)}$. This implies that $\mathbf{w}^{*}$ is flat (Mulayoff and Michaeli, 2020, also see Theorem 1). Now, assume that $\mathbf{w}^{*}$ is flat. By using the fact from Mulayoff and Michaeli (2020, Theorem 2), any flat minimum $\mathbf{w}^{*}$ and $k\in[L]$, it holds that
$\sigma_{\max}(\mathbf{A}_{k})=|m|^{1-\frac{k}{L}}\>\>\text{and}\>\>\sigma_{\max}(\mathbf{B}_{k})=|m|^{\frac{k-1}{L}}.$ (34)
Hence, $\sigma_{\max}(\mathbf{A}_{k})\sigma_{\max}(\mathbf{B}_{k})=|m|^{1-\frac{1}{L}}$ for all $k\in[L]$. ∎
Second, we show that being flat in depth-$2$ matrix factorization is equivalent to being spectral-norm balanced. This result is a direct consequence of Corollary 28 and the AM-GM inequality (see also Josz, 2025, Lemma 5).
Corollary 4.2.
@@
$\mathcal{L}(\mathbf{L},\mathbf{R})=\norm{\mathbf{M}-\mathbf{L}\mathbf{R}^{\top}}_{F}^{2},$ (35)
where $\mathbf{M}\in\mathbb{R}^{d_{L}\times d_{0}}$ is the target matrix and $\mathbf{L}\in\mathbb{R}^{d_{L}\times k}$, $\mathbf{R}\in\mathbb{R}^{d_{0}\times k}$ such that $k\geq\min\{d_{0},d_{L}\}$. Then, $(\mathbf{L}^{*},\mathbf{R}^{*})\in\Omega_{F}$ if and only if
$\sigma_{\max}(\mathbf{L}^{*})=\sigma_{\max}(\mathbf{R}^{*})=\sqrt{\sigma_{\max}(\mathbf{M})}.$ (36)
Remark 4.3.
We discuss further implications of this equivalence in Section 5.
@@
Corollary 4.4.
Consider the optimization objective (18). If $\mathbf{w}^{*}\in\Omega$, then
$\lambda_{\max}\bigl(\nabla^{2}\mathcal{L}(\mathbf{w}^{*})\bigr)\leq 2\sum_{k=1}^{L}\sigma_{\max}(\mathbf{A}_{k})^{2}\sigma_{\max}(\mathbf{B}_{k})^{2},$ (37)
where $\mathbf{A}_{k}=\prod_{i=k+1}^{L}\mathbf{W}_{i}^{*}$ and $\mathbf{B}_{k}=\prod_{i=1}^{k-1}\mathbf{W}_{i}^{*}$ (see Appendix B). This implies that if $\mathbf{w}^{*}$ satisfies $\sigma_{\max}(\mathbf{A}_{k})\sigma_{\max}(\mathbf{B}_{k})=\sigma_{\max}(\mathbf{M})^{1-\frac{1}{L}}$ for all $k\in[L]$, then $\mathbf{w}^{*}$ is flat.
@@
Mulayoff and Michaeli (2020) showed that for any $\mathbf{w}^{*}\in\Omega$,
$\lambda_{\max}\bigl(\nabla^{2}\mathcal{L}(\mathbf{w}^{*})\bigr)\geq 2L\sigma_{\max}(\mathbf{M})^{2(1-1/L)}.$ (38)
Now, assume that $\sigma_{\max}(\mathbf{A}_{k})\sigma_{\max}(\mathbf{B}_{k})=\sigma_{\max}(\mathbf{M})^{1-\frac{1}{L}}$ for all $k\in[L]$. Then by (37),
$\lambda_{\max}\bigl(\nabla^{2}\mathcal{L}(\mathbf{w}^{*})\bigr)=2L\sigma_{\max}(\mathbf{M})^{2(1-1/L)}.$ (39)
This implies that $\mathbf{w}^{*}$ is flat (Mulayoff and Michaeli, 2020, Theorem 1). ∎